determine its molecular formula if its molar mass is 181.313 g mol and it contain 39.7 carbon 2.8 hydrogen 13.4 magnesium and the rest bromine
1
Expert's answer
2016-02-16T00:00:57-0500
The percent of Bromine in the compound is:
100-(39.7+2.8+13.4)=44.1
Assuming that the mass of the compound is 100 g:
m (C)=39.7 g
m (H) = 2.8 g
m (Mg) = 13.4 g
m (Br) = 44.1 g

Considering the molar mass of each component, the amount of moles is:
n (C)=39.7/12.0= 3.3 moles
n (H)=2.8/1.0= 2.8 moles
n (Mg)=13.4/24.3=0.6 moles
n (Br) = 44.1/79.9 = 0.6 moles

Dividing by the lowest, seeking the smallest whole-number ratio:
(C)=3.3/0.6=6
(H)=2.8/0.6=5
(Mg)=0.6/0.6=1
(Br) = 0.6/0.6 = 1

So that the empirical formula of the compound is:
C6H5MgBr, which molar mass is approximately 181.3 g/mol.